A 2.5kg and a 5.0kg ball have an elast collision. Before the collision, the 2.5kg ball was at rest and the other ball had a speed of 3.5m/s. What is the kinetic energy of the 2.5kg ball after the collision?
Answer #1

let m1 = 5.0kg and m2 = 2.5kg

vl/tili = 1,21. <-- A simplification of the kineticenergy equation for elastic collisions during momentumproblems.


vlml m2)=35(50-2.51.167m/s 5.0 + 2.5
v2f = v1f+v1i = 1.167+3.5 = 4.667
m2 v2/ 2 = * 2.5 * 4.6672 = 27 J